# Exploring the Key Features of Repeating Fast Radio Bursts with Machine Learning

### v2

@@ Abstract
Fast radio bursts (FRBs) are enigmatic high-energy events with unknown origins, which are observationally divided into two categories, i.e., repeaters and non-repeaters. However, there are potentially a number of non-repeaters that may be misclassified, as repeating bursts are missed due to the limited sensitivity and observation periods, thus misleading the investigation of their physical properties. In this work, we propose a repeater identification method based on the t-distributed Stochastic Neighbor Embedding (t-SNE) algorithm and apply the classification to the first Canadian Hydrogen Intensity Mapping Experiment Fast Radio Burst (CHIME/FRB) catalog. We find that the spectral morphology parameters, specifically spectral running ($r$), represent the key features for identifying repeaters from the non-repeaters. Also, the results suggest that repeaters are more biased towards narrowband emission, whereas non-repeaters are inclined toward broadband emission. We provide a list of 163 repeater candidates, $5$ of which are confirmed with an updated repeater catalog from CHIME/FRB. Our findings improve our understanding of the various properties underlying repeaters and non-repeaters, as well as guidelines for future FRB detection and categorization.
@@ 1 Introduction
Fast radio bursts (FRBs) are a class of radio flashes with violent explosion energies and millisecond durations. The dispersion measures (DMs) of the FRBs are beyond the maximum contribution of the Milky Way, except for one FRB which has been found to be associated with a magnetar located in the Milky Way (Bochenek et al., 2020). FRBs have received extensive attention since their discovery in 2007 (Lorimer et al., 2007; Thornton et al., 2013; Petroff et al., 2022; Zhang, 2023). Although the physical origin of FRBs is unknown, their unique properties and observations can be utilized to probe fundamental physics, astrophysics, and cosmology (Deng & Zhang, 2014; Macquart et al., 2020; Zhao et al., 2020; Qiu et al., 2022; Yang et al., 2022; Wang & Wei, 2023; Wei & Melia, 2023; Wu et al., 2023; Zhang et al., 2023; Zhao et al., 2023). Numerous experiments, including the Canadian Hydrogen Intensity Mapping Experiment (CHIME) (CHIME/FRB Collaboration et al., 2018), Five-hundred-meter Aperture Spherical radio Telescope (FAST) (Li & Pan, 2016), Tianlai Pathfinder (Yu et al., 2024), and Australian Square Kilometer Array Pathfinder (ASKAP) (Shannon et al., 2024) are currently engaged in FRB research. Their efficient observational capabilities have significantly enriched the FRB sample. So far, over 800 named FRBs have been reported from over 600 sources [^1] https://blinkverse.alkaidos.cn, with more than 200 repeating FRBs from 25 sources and 43 host galaxies located among all observed FRBs (Paine et al., 2024).
@@
In this work, we employ an unsupervised machine learning algorithm, t-distributed Stochastic Neighbour Embedding (t-SNE), with feature pre-selection to identify the potential repeaters in the first CHIME/FRB catalog. The structure of our paper is as follows: In Section 2, we introduce the selected CHIME/FRB data sample, the t-SNE algorithm and corresponding parameter configuration. Section 3 presents the categorization findings of repeaters and non-repeaters found using t-SNE, as well as a detailed comparison of the statistical differences between the two types of FRBs. In Section 4, we summarize our work. The cosmological parameters in this paper are $H_{0}=67.66$ km s-1 Mpc-1, $\Omega_{\rm{b}}=0.04897$, $\Omega_{\rm{m}}=0.3111$, and $\Omega_{\Lambda}=0.6874$ from Planck2018 (Planck Collaboration et al., 2020).
@@ 2.1 Data Sample Selection
We use 536 FRBs published with the first CHIME/FRB catalog, which includes 474 non-repeaters and 62 repeaters. The FRBs in the first CHIME/FRB catalog were detected with observations from July 25th, 2018 to July 1st, 2019. These observations were carried out with a frequency range of $400$ MHz to $800$ MHz (CHIME/FRB Collaboration et al., 2021). We exclude 6 non-repeaters that have no flux or fluence observation data. In addition, some FRBs contain sub-bursts, where a sub-burst refers to a notable local peak in intensity within the profile of an FRB, particularly when this structure appears distinct in frequency from the surrounding emission in the dynamic spectrum, similar to the definition by Sheikh et al. (2024). The FRBs’ sub-bursts are considered as individual bursts in this analysis. So, we compile a sample of 500 non-repeaters and 94 repeaters as original sample (Sorg).
@@ 2.2 FRB Feature Parameters
$S_{\nu}$ (Jy) Peak flux of band-average profile.The value is same for all sub-bursts from a FRB.
@@
$\gamma$ This value represents the spectral shape of every sub-burst. Between each sub-burst from a FRB, the spectral index is different.
@@
$\nu_{\mathrm{Low}}$ (MHz) These values represent the lowest frequency band of pulse detection at a full-width-tenth-maximum. The value between each sub-burst from the FRB is different.
$\Delta t_{\mathrm{BW}}$ (s) Boxcar width means the duration of all sub-pulses combined, including instrumentation, scattering, and redshift broadening effects. The value is same for each sub-burst of the FRB.
$\nu_{\mathrm{High}}$ (MHz) These values represent the highest frequency band of pulse detection at a full-width-tenth-maximum. The value between each subburst from the FRB is different.
@@
Table 1 lists the 11 FRB observational parameters identified by CHIME, including the event name, location, signal-to-noise ratio, and fundamental physical measurements.
@@
Finally, it should be noted that we consider only the lowest frequency ($\nu_{\mathrm{Low}}$) for the frequency characteristics, rather than both the lowest and highest frequency. Observations suggest that the FRB spectrum could span from a few hundred MHz to $\sim 8$ GHz (Gajjar et al., 2018; Pleunis et al., 2021b). The CHIME’s frequency band, i.e. 400400400400800800800800 MHz, covers the lower end of the FRB spectrum. Thus, the FRBs’ highest frequencies are mostly truncated by CHIME’s instrumental upper band limit, and including the highest frequency may limit the intrinsic frequency bandwidth. However, the FRBs’ lowest frequencies are mostly observed within the CHIME frequency band and is more plausible for analyzing CHIME/FRB data. To avoid introducing potential errors for machine learning algorithms, we consider only $\nu_{\mathrm{Low}}$ for characterizing the frequency differences between repeaters and non-repeaters.
@@
t-SNE (t-distributed Stochastic Neighbor Embedding) is a machine-learning algorithm for dimensionality reduction. It is commonly applied to project high-dimensional data into the embedding space with two or three dimensions (van der Maaten & Hinton, 2008; van der Maaten, 2014). The t-SNE algorithm is nonlinear and preserves local data structure by converting high-dimensional distances into conditional probabilities, which guide the arrangement of points in a low-dimensional embedding initialized randomly. This process helps t-SNE reveal clusters and patterns hidden in high-dimensional datasets. Based on random initialization, the t-SNE method can generate different topology structures in the embedding space (Steinhardt et al., 2020). However, in whatever topology structure of the embedding space, data points that were more alike in the original high-dimensional space remain closer together. It should be noted that the coordinate axes of the embedding space have no physical significance. They basically show that closer points are more similar to one another, whereas distant points are less similar. t-SNE has demonstrated robust performance and reliable results when used for dimensionality reduction in classifying gamma-ray bursts (Jespersen et al., 2020; Chen et al., 2024; Zhu et al., 2024). It has also been utilized by Hovis-Afflerbach et al. (2021) to categorize Type-2 Active Galactic Nuclei (AGN) and correct errors in galaxy properties.
@@
The key hyperparameters for the t-SNE initialization process include perplexity, early exaggeration, learning rate, and the number of steps. Perplexity determines the number of nearest neighbors considered when generating conditional probabilities. Larger datasets typically require higher perplexity values to capture more global structures, while with lower perplexity values the algorithm focuses on local structures. The t-SNE documentation suggests that perplexity values within the range of 5-100 generally yield meaningful results [^2] https://distill.pub/2016/misread-tsne/. For optimal outcomes, we recommend considering the stability of the recall value as an auxiliary criterion to determine the most suitable perplexity (see Section 3.3 for details). In this work, perplexity is set to $63.0$ for optimization. Early exaggeration boosts the differences of high-dimensional data points during the initialization, which can help to efficiently separate features in the lower-dimensional embedding. Varying the early exaggeration can significantly affect the topology structure in the embedding space. However, it has little impact on similarity of data points. We use an early exaggeration value of $22.0$. Another crucial hyperparameter is the learning rate. A higher learning rate leads to overly uniform data visualization, where distances between points become nearly equal. Conversely, a lower learning rate tightly compresses data so that discerning the real structure becomes challenging. In our work, the learning rate is finetuned to $265.0$. The detailed description of hyperparameters can be found in Scikit-learn [^3] https://scikit-learn.org/stable/modules/generated/sklearn.manifold.TSNE.html.
@@ 3 RESULTS AND DISCUSSION
The t-SNE dimension reduction results of the Sorg catalog are shown in Figure 2. Significantly, Figure 2 reveals a clear division of FRBs into two distinct clusters, with a pronounced separation between the left and right parts in Figure 2. The red dots represent repeaters, while the green dots represent apparent non-repeaters. The result shows that almost all observed repeaters are gathered on the right side, while the left side is almost exclusively populated by observed apparent non-repeaters. As discussed in Section 2.3, the coordinates of the t-SNE embedding space represent no physical meaning. However, in the embedding space, FRBs close to each other have higher similarities. There is a significant fraction of the apparent non-repeaters located on the right side of Figure 2, which are mixed with the observationally defined repeaters. This suggests that such non-repeating FRBs could be candidates for repeating FRBs. The fact that these FRBs are currently classified as non-repeaters may be due to observational limitations. This also implies a considerable distinction in the characteristics of repeaters
@@
We further categorize the FRBs as repeaters and non-repeaters in the t-SNE embedding space via the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithm (McInnes et al., 2017). The HDBSCAN algorithm, which is an extension of DBlCAN (Density-Based Spatial Clustering of Applications with Noise), is a density-based clustering algorithm designed to identify density-connected regions by calculating distances between data points. The HDBSCAN algorithm introduces a hierarchical structure by constructing a distance matrix and generating a directed weighted graph, which it traverses to assign data points to clusters. HDBSCAN exhibits a more powerful clustering performance and is capable of discovering clusters of arbitrary shape and size. The clustering results are shown in Figure 3, where the FRB samples are divided into two clusters. One of the two clusters, which contains a significant proportion of repeaters, is labeled as the repeater cluster (more than $10\%$ of the FRBs are repeaters), while another cluster is termed as the non-repeater cluster. Although the rest FRBs in the repeater cluster have only one burst detected, their features are close to the repeating FRBs in the selected parameter space. We label such FRBs as the repeater candidates. Table 3 lists the detailed content of the two clusters. On the other hand, there is one repeating FRB categorized in the non-repeater cluster. We further discuss the results verification in Section 3.4.
@@
We further explore potential correlations between the embedding space clustering results and FRB morphology as well as host galaxy properties. Based on Pleunis et al. (2021a), we randomly select several FRBs from the first CHIME/FRB catalog and categorize them as simple broadband bursts, simple narrowband bursts, complex broadband bursts with multiple peaks, complex narrowband bursts with multiple peaks, and downward drifting bursts for comparison with Figure 3. The results indicate that broadband bursts consistently belong to the non-repeater cluster, while narrowband bursts are always classified within the repeater cluster. However, there is no clear correlation between the two clusters and complex, simple, or downward drifting bursts. Host galaxy properties are also crucial for understanding the origins and characteristics of FRBs. According to the recent statistical results by Ma et al. (2024), nine FRBs in the first CHIME/FRB catalog are localized, including five repeaters and four non-repeaters. We compare the host galaxy morphological properties and stellar masses of these nine FRBs with the t-SNE dimensionality reduction results. The comparison indicates that there is no significant difference in host galaxy properties between repeaters and non-repeaters, consistent with the conclusions of Ma et al. (2024).
@@
In the following sections, we refer to the FRBs catalog sample categorized in the t-SNE embedding space as the grouped sample (Sgrp), which consist of 94 repeaters, 163 repeater candidates, and 337 non-repeaters.
@@ 3.2 Feature Importance
In the previous section, we show the results of the unsupervised FRB classification in the t-SNE embedding space based on seven observation parameters. To gain a deeper understanding of the FRBs classification process, we explore which observable parameters of FRBs are more significant in machine learning classification. For better visualization, we apply color-coding for each FRB of the two clusters based on the corresponding values of $r$, $\gamma$, $S_{\nu}$, $F_{\nu}$, $\Delta t_{\mathrm{ST}}$, $\Delta t_{\mathrm{WS}}$, and $\nu_{\mathrm{Low}}$, respectively. As shown in Figure 4, the values of $r$, $\gamma$, and $\nu_{\mathrm{Low}}$ show obvious gradients in the t-SNE embedding space. Other physical parameters exhibit no noticeable regular variation between these two clusters. By visual inspection, t-SNE has roughly approximated the FRBs into two clusters according to $r$ and $\gamma$. However, there is no clear boundary for $r$ and $\gamma$ alone between the two clusters. Specifically, the $r$ values of some FRBs in the repeater cluster are greater than some non-repeater clusters’.
Figure 4: t-SNE embedding space, coloured based on $r$, $\gamma$, $S_{\nu}$, $F_{\nu}$, $\Delta t_{\mathrm{ST}}$, $\Delta t_{\mathrm{WS}}$, and $\nu_{\mathrm{Low}}$.
@@
${\rm Recall}=\frac{\rm T_{P}}{\rm T_{P}+\rm F_{N}},$ (1)
where TP, i.e. true positive, denotes the number of repeaters that fall in the repeater clusters, and FN, i.e. false negative, denotes the number of repeaters that fall in the non-repeater clusters. We adopt a model validation technique named permutation feature importance (Altmann et al., 2010; Yang et al., 2023). Specifically, we sequentially shuffle the values of the seven FRB observed parameters and apply the value-shuffled data to the same t-SNE classification process. By shuffling the values of one particular FRB observation parameter, the similarity information carried by such parameters is disrupted, leading to a loss in recall. The permutation feature importance is quantified by the recall loss rate caused:
@@
Figure 5 illustrates the permutation feature importance, i.e. the recall loss rate of the seven FRB observational parameters. The results show that the parameters of $r$ and $\gamma$ are the most important, i.e. the loss caused by the permutation of $r$ reaches $22.98\%$ and the second-ranking $\gamma$ causes a $2.77\%$ loss. The importance of the other observational parameters shows relatively minor differences. Therefore, we show that in the unsupervised classification of FRBs, the parameter $r$ plays a pivotal role in differentiating repeaters and non-repeaters.

### v3

@@ Abstract
Fast radio bursts (FRBs) are enigmatic high-energy events with unknown origins, which are observationally divided into two categories, i.e., repeaters and non-repeaters. However, there are potentially a number of non-repeaters that may be misclassified, as repeating bursts are missed due to the limited sensitivity and observation periods, thus misleading the investigation of their physical properties. In this work, we propose a repeater identification method based on the t-distributed Stochastic Neighbor Embedding algorithm and apply the classification to the first Canadian Hydrogen Intensity Mapping Experiment Fast Radio Burst (CHIME/FRB) catalog. We find that the spectral morphology parameters, specifically spectral running ($r$), represent the key features for identifying repeaters from the non-repeaters. Also, the results suggest that repeaters are more biased toward narrowband emission, whereas non-repeaters are inclined toward broadband emission. We provide a list of 163 repeater candidates, five of which are confirmed with an updated repeater catalog from CHIME/FRB. Our findings improve our understanding of the various properties underlying repeaters and non-repeaters, as well as guidelines for future FRB detection and categorization.
@@ 1 Introduction
Fast radio bursts (FRBs) are a class of radio flashes with violent explosion energies and millisecond durations. The dispersion measures (DMs) of the FRBs are beyond the maximum contribution of the Milky Way, except for one FRB, which has been found to be associated with a magnetar located in the Milky Way (Bochenek et al., 2020). FRBs have received extensive attention since their discovery in 2007 (Lorimer et al., 2007; Thornton et al., 2013; Petroff et al., 2022; Zhang, 2023). Although the physical origin of FRBs is unknown, their unique properties and observations can be utilized to probe fundamental physics, astrophysics, and cosmology (Deng & Zhang, 2014; Macquart et al., 2020; Zhao et al., 2020; Qiu et al., 2022; Yang et al., 2022; Wang & Wei, 2023; Wei & Melia, 2023; Wu et al., 2023; Zhang et al., 2023; Zhao et al., 2023). Numerous experiments, including the Canadian Hydrogen Intensity Mapping Experiment (CHIME) (CHIME/FRB Collaboration et al., 2018), Five-hundred-meter Aperture Spherical radio Telescope (Li & Pan, 2016), Tianlai Pathfinder (Yu et al., 2024), and Australian Square Kilometer Array Pathfinder (Shannon et al., 2024) are currently engaged in FRB research. Their efficient observational capabilities have significantly enriched the FRB sample. So far, over 800 named FRBs have been reported from over 600 sources [^1] https://blinkverse.alkaidos.cn, with more than 200 repeating FRBs from 25 sources and 43 host galaxies located among all observed FRBs (Paine et al., 2024).
@@
In this work, we employ an unsupervised machine learning algorithm, t-distributed Stochastic Neighbour Embedding (t-SNE), with feature preselection to identify the potential repeaters in the first CHIME/FRB catalog. The structure of our paper is as follows: In Section 2, we introduce the selected CHIME/FRB data sample, the t-SNE algorithm and corresponding parameter configuration. Section 3 presents the categorization findings of repeaters and non-repeaters found using t-SNE, as well as a detailed comparison of the statistical differences between the two types of FRBs. In Section 4, we summarize our work. The cosmological parameters in this paper are $H_{0}=67.66$ km s-1 Mpc-1, $\Omega_{\rm{b}}=0.04897$, $\Omega_{\rm{m}}=0.3111$, and $\Omega_{\Lambda}=0.6874$ from Planck2018 (Planck Collaboration et al., 2020).
@@ 2.1 Data Sample Selection
We use 536 FRBs published with the first CHIME/FRB catalog, which includes 474 non-repeaters and 62 repeaters. The FRBs in the first CHIME/FRB catalog were detected with observations from 2018 July 25 to 2019 July 1. These observations were carried out with a frequency range of $400$ MHz to $800$ MHz (CHIME/FRB Collaboration et al., 2021). We exclude six non-repeaters that have no flux or fluence observation data. In addition, some FRBs contain sub-bursts, where a sub-burst refers to a notable local peak in intensity within the profile of an FRB, particularly when this structure appears distinct in frequency from the surrounding emission in the dynamic spectrum, similar to the definition by Sheikh et al. (2024). The FRBs’ sub-bursts are considered as individual bursts in this analysis. So, we compile a sample of 500 non-repeaters and 94 repeaters as original sample (Sorg).
@@ 2.2 FRB Feature Parameters
$S_{\nu}$ (Jy) Peak flux of band-average profile.The value is the same for all sub-bursts from an FRB.
@@
$\gamma$ This value represents the spectral shape of every sub-burst. Between each sub-burst from an FRB, the spectral index is different.
@@
$\nu_{\mathrm{Low}}$ (MHz) These values represent the lowest frequency band of pulse detection at a full width tenth maximum. The value between each sub-burst from the FRB is different.
$\Delta t_{\mathrm{BW}}$ (s) Boxcar width means the duration of all sub-pulses combined, including instrumentation, scattering, and redshift broadening effects. The value is the same for each sub-burst of the FRB.
$\nu_{\mathrm{High}}$ (MHz) These values represent the highest frequency band of pulse detection at a full-width-tenth-maximum. The value between each sub-burst from the FRB is different.
@@
Table 1 lists the 11 FRB observational parameters identified by CHIME, including the event name, location, signal-to-noise ratio (SNR), and fundamental physical measurements.
@@
Finally, it should be noted that we consider only the lowest frequency ($\nu_{\mathrm{Low}}$) for the frequency characteristics, rather than both the lowest and highest frequency. Observations suggest that the FRB spectrum could span from a few hundred MHz to $\sim 8$ GHz (Gajjar et al., 2018; Pleunis et al., 2021b). The CHIME’s frequency band, i.e. 400400400400800800800800 MHz, covers the lower end of the FRB spectrum. Thus, the FRBs’ highest frequencies are mostly truncated by CHIME’s instrumental upper band limit, and including the highest frequency may limit the intrinsic frequency bandwidth. However, the FRBs’ lowest frequencies are mostly observed within the CHIME frequency band and are more plausible for analyzing CHIME/FRB data. To avoid introducing potential errors for machine learning algorithms, we consider only $\nu_{\mathrm{Low}}$ for characterizing the frequency differences between repeaters and non-repeaters.
@@
t-SNE is a machine learning algorithm for dimensionality reduction. It is commonly applied to project high-dimensional data into the embedding space with two or three dimensions (van der Maaten & Hinton, 2008; van der Maaten, 2014). The t-SNE algorithm is nonlinear and preserves local data structure by converting high-dimensional distances into conditional probabilities, which guide the arrangement of points in a low-dimensional embedding initialized randomly. This process helps t-SNE reveal clusters and patterns hidden in high-dimensional datasets. Based on random initialization, the t-SNE method can generate different topology structures in the embedding space (Steinhardt et al., 2020). However, in whatever topology structure of the embedding space, data points that were more alike in the original high-dimensional space remain closer together. It should be noted that the coordinate axes of the embedding space have no physical significance. They basically show that closer points are more similar to one another, whereas distant points are less similar. t-SNE has demonstrated robust performance and reliable results when used for dimensionality reduction in classifying gamma-ray bursts (Jespersen et al., 2020; Chen et al., 2024; Zhu et al., 2024). It has also been utilized by Hovis-Afflerbach et al. (2021) to categorize Type 2 active galactic nuclei and correct errors in galaxy properties.
@@
The key hyperparameters for the t-SNE initialization process include perplexity, early exaggeration, learning rate, and the number of steps. Perplexity determines the number of nearest neighbors considered when generating conditional probabilities. Larger datasets typically require higher perplexity values to capture more global structures, while with lower perplexity values the algorithm focuses on local structures. The t-SNE documentation suggests that perplexity values within the range of 5-100 generally yield meaningful results [^2] https://distill.pub/2016/misread-tsne/. For optimal outcomes, we recommend considering the stability of the recall value as an auxiliary criterion to determine the most suitable perplexity (see Section 3.3 for details). In this work, perplexity is set to $63.0$ for optimization. Early exaggeration boosts the differences of high-dimensional data points during the initialization, which can help to efficiently separate features in the lower-dimensional embedding. Varying the early exaggeration can significantly affect the topology structure in the embedding space. However, it has little impact on similarity of data points. We use an early exaggeration value of $22.0$. Another crucial hyperparameter is the learning rate. A higher learning rate leads to overly uniform data visualization, where distances between points become nearly equal. Conversely, a lower learning rate tightly compresses data so that discerning the real structure becomes challenging. In our work, the learning rate is fine-tuned to $265.0$. A detailed description of hyperparameters can be found in Scikit-learn [^3] https://scikit-learn.org/stable/modules/generated/sklearn.manifold.TSNE.html.
@@ 3 RESULTS AND DISCUSSION
The t-SNE dimension reduction results of the Sorg catalog are shown in Figure 2. Significantly, Figure 2 reveals a clear division of FRBs into two distinct clusters, with a pronounced separation between the left and right parts in Figure 2. The red dots represent repeaters, while the green dots represent apparent non-repeaters. The result shows that almost all observed repeaters are gathered on the right side, while the left side is almost exclusively populated by observed apparent non-repeaters. As discussed in Section 2.3, the coordinates of the t-SNE embedding space represent no physical meaning. However, in the embedding space, FRBs close to each other have higher similarities. A significant fraction of the apparent non-repeaters located on the right side of Figure 2, which are mixed with the observationally defined repeaters. This suggests that such non-repeating FRBs could be candidates for repeating FRBs. The fact that these FRBs are currently classified as non-repeaters may be due to observational limitations. This also implies a considerable distinction in the characteristics of repeaters
@@
We further categorize the FRBs as repeaters and non-repeaters in the t-SNE embedding space via the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithm (McInnes et al., 2017). The HDBSCAN algorithm, which is an extension of Density-Based Spatial Clustering of Applications with Noise, is a density-based clustering algorithm designed to identify density-connected regions by calculating distances between data points. The HDBSCAN algorithm introduces a hierarchical structure by constructing a distance matrix and generating a directed weighted graph, which it traverses to assign data points to clusters. HDBSCAN exhibits a more powerful clustering performance and is capable of discovering clusters of arbitrary shape and size. The clustering results are shown in Figure 3, where the FRB samples are divided into two clusters. One of the two clusters, which contains a significant proportion of repeaters, is labeled as the repeater cluster (more than $10\%$ of the FRBs are repeaters), while another cluster is termed as the non-repeater cluster. Although the rest of the FRBs in the repeater cluster have only one burst detected, their features are close to the repeating FRBs in the selected parameter space. We label such FRBs as the repeater candidates. Table 3 lists the detailed content of the two clusters. On the other hand, there is one repeating FRB categorized in the non-repeater cluster. We further discuss the verification of results in Section 3.4.
@@
We further explore potential correlations between the embedding space clustering results and FRB morphology as well as host galaxy properties. Based on Pleunis et al. (2021a), we randomly select several FRBs from the first CHIME/FRB catalog and categorize them as simple broadband bursts, simple narrowband bursts, complex broadband bursts with multiple peaks, complex narrowband bursts with multiple peaks, and downward-drifting bursts for comparison with Figure 3. The results indicate that broadband bursts consistently belong to the non-repeater cluster, while narrowband bursts are always classified within the repeater cluster. However, there is no clear correlation between the two clusters and complex, simple, or downward-drifting bursts. Host galaxy properties are also crucial for understanding the origins and characteristics of FRBs. According to the recent statistical results by Ma et al. (2024), nine FRBs in the first CHIME/FRB catalog are localized, including five repeaters and four non-repeaters. We compare the host galaxy morphological properties and stellar masses of these nine FRBs with the t-SNE dimensionality reduction results. The comparison indicates that there is no significant difference in host galaxy properties between repeaters and non-repeaters, consistent with the conclusions of Ma et al. (2024).
@@
In the following sections, we refer to the FRB catalog sample categorized in the t-SNE embedding space as the grouped sample (Sgrp), which consists of 94 repeaters, 163 repeater candidates, and 337 non-repeaters.
@@ 3.2 Feature Importance
In the previous section, we show the results of the unsupervised FRB classification in the t-SNE embedding space based on seven observation parameters. To gain a deeper understanding of the FRB classification process, we explore which observable parameters of FRBs are more significant in machine learning classification. For better visualization, we apply color coding for each FRB of the two clusters based on the corresponding values of $r$, $\gamma$, $S_{\nu}$, $F_{\nu}$, $\Delta t_{\mathrm{ST}}$, $\Delta t_{\mathrm{WS}}$, and $\nu_{\mathrm{Low}}$, respectively. As shown in Figure 4, the values of $r$, $\gamma$, and $\nu_{\mathrm{Low}}$ show obvious gradients in the t-SNE embedding space. Other physical parameters exhibit no noticeable regular variation between these two clusters. By visual inspection, t-SNE has roughly approximated the FRBs into two clusters according to $r$ and $\gamma$. However, there is no clear boundary for $r$ and $\gamma$ alone between the two clusters. Specifically, the $r$ values of some FRBs in the repeater cluster are greater than some non-repeater clusters’.
Figure 4: t-SNE embedding space, colored based on $r$, $\gamma$, $S_{\nu}$, $F_{\nu}$, $\Delta t_{\mathrm{ST}}$, $\Delta t_{\mathrm{WS}}$, and $\nu_{\mathrm{Low}}$.
@@
${\rm Recall}=\frac{T_{\rm P}}{T_{\rm P}+F_{\rm N}},$ (1)
where $T_{\rm P}$, i.e. true positive, denotes the number of repeaters that fall in the repeater clusters, and $F_{\rm N}$, i.e. false negative, denotes the number of repeaters that fall in the non-repeater clusters. We adopt a model validation technique named permutation feature importance (Altmann et al., 2010; Yang et al., 2023). Specifically, we sequentially shuffle the values of the seven FRB observed parameters and apply the value-shuffled data to the same t-SNE classification process. By shuffling the values of one particular FRB observation parameter, the similarity information carried by such parameters is disrupted, leading to a loss in recall. The permutation feature importance is quantified by the recall loss rate caused:
@@
Figure 5 illustrates the importance of the permutation feature, i.e. the recall loss rate of the seven FRB observational parameters. The results show that the parameters of $r$ and $\gamma$ are the most important, i.e. the loss caused by the permutation of $r$ reaches $22.98\%$ and the second-ranking $\gamma$ causes a $2.77\%$ loss. The importance of the other observational parameters shows relatively minor differences. Therefore, we show that in the unsupervised classification of FRBs, the parameter $r$ plays a pivotal role in differentiating repeaters and non-repeaters.